OpenStudy (anonymous):

Evaluate each function for the domain {1, 2, 3} f(x)=4^x y=40(1/2)^x Graph each function f(x)=2.5^x f(x)=1/2 times 3^x

OpenStudy (anonymous):

Will medal and fan!!

OpenStudy (jdoe0001):

\(\qquad \qquad \qquad x=\{1,2,3\}\\ \begin{array}{llll} \qquad &f(x)=4^x\qquad &y=40\left( \frac{1}{2} \right)^x\\ \\\hline\\ 1&f(x)=4^1&y=40\left( \frac{1}{2} \right)^1\\ 2&f(x)=4^2&y=40\left( \frac{1}{2} \right)^2\\ 3&f(x)=4^3&y=40\left( \frac{1}{2} \right)^3 \end{array}\)

OpenStudy (jdoe0001):

for the graphing, make a table of values, or use a graphing calculator if you're expected to do so

OpenStudy (anonymous):

ohh thank you so much